# Single-stream Policy Optimization

### v1

@@
Abstract: We revisit policy-gradient optimization for Large Language Models (LLMs) from a single-stream perspective. Prevailing group-based methods like GRPO reduce variance with on-the-fly baselines but suffer from critical flaws: frequent degenerate groups erase learning signals, and synchronization barriers hinder scalability. We introduce Single-stream Policy Optimization (SPO), which eliminates these issues by design. SPO replaces per-group baselines with a persistent, KL-adaptive value tracker and normalizes advantages globally across the batch, providing a stable, low-variance learning signal for every sample. Being group-free, SPO enables higher throughput and scales effectively in long-horizon or tool-integrated settings where generation times vary. Furthermore, the persistent value tracker naturally enables an adaptive curriculum via prioritized sampling. Experiments using Qwen3-8B show that SPO converges more smoothly and attains higher accuracy than GRPO, while eliminating computation wasted on degenerate groups. Ablation studies confirm that SPO’s gains stem from its principled approach to baseline estimation and advantage normalization, offering a more robust and efficient path for LLM reasoning. Across five hard math benchmarks with Qwen3-8B, SPO improves the average maj@32 by $3.4\leavevmode\nobreak\ \text{percentage points}\leavevmode\nobreak\ (\mathrm{pp})$ over GRPO, driven by substantial absolute point gains on challenging datasets, including $7.3\leavevmode\nobreak\ \mathrm{pp}$ on BRUMO 25, $4.4\leavevmode\nobreak\ \mathrm{pp}$ on AIME 25, $3.3\leavevmode\nobreak\ \mathrm{pp}$ on HMMT 25, and achieves consistent relative gain in pass@$k$ across the evaluated $k$ values. SPO’s success challenges the prevailing trend of adding incidental complexity to RL algorithms, highlighting a path where fundamental principles, not architectural workarounds, drive the next wave of progress in LLM reasoning.
@@ 1 Introduction
The benefits of this principled approach are clear: SPO is inherently more scalable and eliminates the computational waste of degenerate groups. Our experiments confirm these advantages, demonstrating that SPO consistently outperforms GRPO on challenging reasoning benchmarks, improving the absolute point gains on challenging datasets, including $7.3\leavevmode\nobreak\ \text{percentage points}\leavevmode\nobreak\ (\mathrm{pp})$ on BRUMO 25, $4.4\leavevmode\nobreak\ \mathrm{pp}$ on AIME 25, $3.3\leavevmode\nobreak\ \mathrm{pp}$ on HMMT 25, and the pass@$k$ curves of SPO are above GRPO for all $k$s. The scalability benefit is particularly pronounced in agentic settings. Our simulations, designed to model these variable-time scenarios, show that SPO’s group-free design can achieve a $4.35\times$ training throughput speedup by eliminating group synchronization bottlenecks. SPO thus provides a more robust foundation for modern LLM optimization, prompting a re-evaluation of essential versus incidental complexity in the field.
@@ 2 Related Work
Several other works have proposed improvements to group-based methods. Zheng et al. [43] introduce GRESO, an online filtering algorithm that leverages reward training dynamics to predict and skip uninformative prompts before generation, significantly reducing rollout overhead. Liu et al. [22] propose “Lite PPO”, which simplifies RLVR training to only advantage normalization and token-level loss aggregation.
@@ 4.1 A KL-Adaptive Value Tracker
$\alpha(x)=\rho(x)\alpha_{-1}(x)+r(x,y),\qquad\beta(x)=\rho(x)\beta_{-1}(x)+(1-r(x,y)),\qquad\hat{v}(x)=\frac{\alpha(x)}{\alpha(x)+\beta(x)}.$ (5)
@@ 4.2 Advantage Estimation and Policy Optimization
Using the baseline from the previous step ensures that it is independent of the action taken at step $i$, preserving the unbiasedness of the policy gradient estimate. Since $v_{-1}(x)$ is independent of $y\sim\pi_{\theta_{i}}(\cdot|x)$, $\mathbb{E}\!\left[(r-v_{i-1}(x))\,\nabla_{\theta}\log\pi\right]=\nabla J(\theta)$ [37]. Instead of normalizing advantages on a per-prompt basis in a group [27, 31], SPO normalizes them across an entire batch of prompts $\mathcal{B}$ [17, 25, 3, 20]. The normalized advantage $\tilde{A}(x,y)$ is computed as:
@@ 5.2 Empirical Comparison with GRPO
Table 1: Comparison of GRPO, and SPO on five benchmarks using maj@32 and avg@32. Averages are shown in the last column. Bold indicates the better-performing method for each metric.
@@
Our experiments demonstrate that SPO outperforms the GRPO baseline on aggregate metrics when training the Qwen-8B model. As shown in Table 1, SPO achieves superior weighted average scores on both primary metrics. It obtains a maj@32 of $63.8$ compared to GRPO’s $60.4$, a significant improvement of $3.4\leavevmode\nobreak\ \text{percentage points}\leavevmode\nobreak\ (\mathrm{pp})$. This aggregate strength is driven by remarkable consistency, as SPO outperforms GRPO on the maj@32 metric across all five benchmarks. The performance gap is most pronounced on BRUMO 25, where SPO achieves a substantial $7.3\leavevmode\nobreak\ \mathrm{pp}$ ($64.0$ vs. $56.7$). Further significant gains are seen on AIME 25 ($4.4\leavevmode\nobreak\ \mathrm{pp}$) and HMMT 25 ($3.3\leavevmode\nobreak\ \mathrm{pp}$ points), underscoring the robustness of SPO’s improvements. While GRPO remains competitive on the avg@32 metric in some cases, SPO’s consistent and significant advantage in maj@32 suggests it learns more robust and repeatable solutions, a key goal for reliable reasoning models.
These findings are mirrored in the pass@$k$ performance shown in Figure 2. The weighted average curve (Figure 2(f)) shows a clear and consistent advantage for SPO across all values of $k$, translating to an average improvement of approximately $2.4\leavevmode\nobreak\ \mathrm{pp}$. While the performance on avg@32 is more competitive on a per-benchmark basis, SPO’s strong overall performance underscores the stability and effectiveness of its learning signal.
@@ 5.3 Analysis of Signal Efficiency and Stability
Figure 3: Signal Efficiency and Stability Analysis of SPO vs. GRPO. (a) The majority of GRPO samples fall into “degenerate groups” (blue line), yielding zero advantage and no learning signal. In contrast, SPO’s rate of near-zero advantages (red/green lines) increases as its value tracker learns, indicating successful learning, not wasted computation. (b) Comparison of unnormalized advantage variance. SPO’s baseline (red) provides a stable, low-variance signal, significantly reducing the variance of the raw reward (denoted as “SPO No Baseline” as green). GRPO’s effective advantage (blue), calculated only on its non-degenerate samples, is highly volatile, even more so than the raw reward variance (green line, labeled “SPO No Baseline”), revealing the instability of its on-the-fly approach.
@@
Signal Efficiency and Information Loss. Figure 3(a) reports the fraction of ineffective samples. For GRPO (blue), the share of samples in degenerate groups rises from roughly 60% to over 80%, yielding zero advantage and no gradient. For SPO, we instead track the proportion of near-zero advantages under two diagnostic tolerances: exact zero (threshold=0.0; red) and a small band $\lvert A\rvert<0.02$ (threshold=0.02; green). Exact zeros remain rare throughout training (red near zero), while the $\lvert A\rvert<0.02$ share (green) gradually increases as the value tracker $\hat{v}$ becomes more accurate and residuals shrink on mastered prompts. This trend is expected and desirable: it reflects accurate prediction rather than signal loss. Unlike GRPO’s degenerate groups, these SPO samples are not discarded—they still produce well-defined gradients and contribute to learning. Notably, even under the stricter 0.02 tolerance, SPO’s near-zero ratio remains far below GRPO’s degenerate rate, underscoring SPO’s efficient use of compute.
@@ Appendix A SPO Initialization
We show the SPO initialization procedure in Algorithm 2. In the experiments, we use $n_{0}=8$ to have a good estimation of initial baseline tracker. We ablate the setting where we use no offline estimation and rely on the online moving estimator. Our early-stage experiments show that without offline estimation the training is unstable in the first steps.
@@
Practically, one may concern about the extra cost during the offline estimation of $\hat{v}_{0}$. We note that we share the offline estimation for our experiments [^4] https://huggingface.co/datasets/dingzihan737/SPO_Qwen3-8B_DAPO_16k_Tool so that people could skip this process and directly load our datasets, and there are datasets like Polaris [2] that pre-compute accuracy for Deepseek-R1-Distill-Qwen-7B [31]. The cost can be amortized across the experiments people run themselves, and we will share more (dataset, base_model) combinations to facilitate experiment efficiency.

### v2

@@
Abstract: We revisit policy-gradient optimization for Large Language Models (LLMs) from a single-stream perspective. Prevailing group-based methods like GRPO reduce variance with on-the-fly baselines but suffer from critical flaws: frequent degenerate groups erase learning signals, and synchronization barriers hinder scalability. We introduce Single-stream Policy Optimization (SPO), which eliminates these issues by design. SPO replaces per-group baselines with a persistent, KL-adaptive value tracker and normalizes advantages globally across the batch, providing a stable, low-variance learning signal for every sample. Being group-free, SPO enables higher throughput and scales effectively in long-horizon or tool-integrated settings where generation times vary. Furthermore, the persistent value tracker naturally enables an adaptive curriculum via prioritized sampling. Experiments using Qwen3-8B show that SPO converges more smoothly and attains higher accuracy than GRPO, while eliminating computation wasted on degenerate groups. Ablation studies confirm that SPO’s gains stem from its principled approach to baseline estimation and advantage normalization, offering a more robust and efficient path for LLM reasoning. Across five hard math benchmarks with Qwen3-8B, SPO improves the average maj@32 by $3.4\penalty 10000\ \text{percentage points}\penalty 10000\ (\mathrm{pp})$ over GRPO, driven by substantial absolute point gains on challenging datasets, including $7.3\penalty 10000\ \mathrm{pp}$ on BRUMO 25, $4.4\penalty 10000\ \mathrm{pp}$ on AIME 25, $3.3\penalty 10000\ \mathrm{pp}$ on HMMT 25, and achieves consistent relative gain in pass@$k$ across the evaluated $k$ values. SPO’s success challenges the prevailing trend of adding incidental complexity to RL algorithms, highlighting a path where fundamental principles, not architectural workarounds, drive the next wave of progress in LLM reasoning.
@@ 1 Introduction
The benefits of this principled approach are clear: SPO is inherently more scalable and eliminates the computational waste of degenerate groups. Our experiments confirm these advantages, demonstrating that SPO consistently outperforms GRPO on challenging reasoning benchmarks, improving the absolute point gains on challenging datasets, including $7.3\penalty 10000\ \text{percentage points}\penalty 10000\ (\mathrm{pp})$ on BRUMO 25, $4.4\penalty 10000\ \mathrm{pp}$ on AIME 25, $3.3\penalty 10000\ \mathrm{pp}$ on HMMT 25, and the pass@$k$ curves of SPO are above GRPO for all $k$s. The scalability benefit is particularly pronounced in agentic settings. Our simulations, designed to model these variable-time scenarios, show that SPO’s group-free design can achieve a $4.35\times$ training throughput speedup by eliminating group synchronization bottlenecks. SPO thus provides a more robust foundation for modern LLM optimization, prompting a re-evaluation of essential versus incidental complexity in the field.
@@ 2 Related Work
Several other works have proposed improvements to group-based methods. Zheng et al. [44] introduce GRESO, an online filtering algorithm that leverages reward training dynamics to predict and skip uninformative prompts before generation. Qu et al. [28] introduce a Bayesian estimation of the prompt accuracy and use it to form a bandit strategy, significantly reducing rollout overhead. Liu et al. [25] propose “Lite PPO”, which simplifies RLVR training to only advantage normalization and token-level loss aggregation.
@@ 4.1 A KL-Adaptive Value Tracker
$\alpha(x)=\rho(x)\alpha_{-1}(x)+r(x,y),\quad\beta(x)=\rho(x)\beta_{-1}(x)+(1-r(x,y)),\quad\hat{v}(x)=\frac{\alpha(x)}{\alpha(x)+\beta(x)}.$ (5)
@@ 4.2 Advantage Estimation and Policy Optimization
Using the baseline from the previous step ensures that it is independent of the action taken at step $i$, preserving the unbiasedness of the policy gradient estimate. While the reward $r(x,y)$ is typically a direct outcome signal, SPO’s framework is also compatible with more sophisticated reward functions. For instance, recent work like InfAlign [4] demonstrates how to calibrate and transform the reward signal to be “inference-aware,” directly optimizing for procedures like Best-of-$N$ sampling. Such transformed rewards can be seamlessly integrated into SPO by replacing the standard $r(x,y)$ in the advantage calculation. Since $v_{-1}(x)$ is independent of $y\sim\pi_{\theta_{i}}(\cdot|x)$, $\mathbb{E}\!\left[(r-v_{i-1}(x))\,\nabla_{\theta}\log\pi\right]=\nabla J(\theta)$ [39]. Instead of normalizing advantages on a per-prompt basis in a group [34, 11], SPO normalizes them across an entire batch of prompts $\mathcal{B}$ [19, 31, 3, 23]. The normalized advantage $\tilde{A}(x,y)$ is computed as:
@@ 5.2 Empirical Comparison with GRPO
Table 1: Comparison of GRPO and SPO on five benchmarks using maj@32 and avg@32. Averages are shown in the last column. Bold indicates the better-performing method for each metric.
@@
Our experiments demonstrate that SPO outperforms the GRPO baseline on aggregate metrics when training the Qwen-8B model. As shown in Table 1, SPO achieves superior weighted average scores on both primary metrics. It obtains a maj@32 of $63.8$ compared to GRPO’s $60.4$, a significant improvement of $3.4\penalty 10000\ \text{percentage points}\penalty 10000\ (\mathrm{pp})$. This aggregate strength is driven by remarkable consistency, as SPO outperforms GRPO on the maj@32 metric across all five benchmarks. The performance gap is most pronounced on BRUMO 25, where SPO achieves a substantial $7.3\penalty 10000\ \mathrm{pp}$ ($64.0$ vs. $56.7$). Further significant gains are seen on AIME 25 ($4.4\penalty 10000\ \mathrm{pp}$) and HMMT 25 ($3.3\penalty 10000\ \mathrm{pp}$ points), underscoring the robustness of SPO’s improvements. Notably, these benchmarks have minimal data contamination [5], allowing them to serve as a true test of generalization. This demonstrates that our SPO method improves the model’s ability to generalize rather than simply overfit to the training data, a risk exemplified by the DAPO dataset’s strong correlation with AIME 24. While GRPO remains competitive on the avg@32 metric in some cases, SPO’s consistent and significant advantage in maj@32 suggests it learns more robust and repeatable solutions, a key goal for reliable reasoning models.
These findings are mirrored in the pass@$k$ performance shown in Figure 2. The weighted average curve (Figure 2(f)) shows a clear and consistent advantage for SPO across all values of $k$, translating to an average improvement of approximately $2.4\penalty 10000\ \mathrm{pp}$. While the performance on avg@32 is more competitive on a per-benchmark basis, SPO’s strong overall performance underscores the stability and effectiveness of its learning signal. We provide additional ablation studies on $A^{*}$-PO, SPO with no baseline, and SPO with no offline initialization in Appendix E.
@@ 5.3 Analysis of Signal Efficiency and Stability
Figure 3: Signal Efficiency and Stability Analysis of SPO vs. GRPO. (a) GRPO suffers from a high ratio of degenerate groups (blue), which yield no learning signal. In contrast, SPO’s rate of near-zero advantages (red/green) increases as the model learns, reflecting prediction accuracy rather than wasted computation. (b) SPO’s baseline (red) provides a stable, low-variance signal, significantly reducing the raw reward variance (green). GRPO’s effective advantage (blue), calculated only on non-degenerate samples, is highly volatile and unstable.
@@
Signal Efficiency and Information Loss. Figure 3(a) reports the fraction of ineffective samples. For GRPO (blue), the share of samples in degenerate groups rises from roughly 60% to over 80%, yielding zero advantage and no gradient. For SPO, we instead track the proportion of near-zero advantages under two diagnostic tolerances, $\lvert A\rvert\leq\tau$, with values of $\tau=10^{-4}$ (red) and $\tau=0.02$ (green). Advantages under the tight tolerance $\tau=10^{-4}$ remain rare throughout training (red line), while the $\lvert A\rvert\leq 0.02$ share (green) gradually increases as the value tracker $\hat{v}$ becomes more accurate and residuals shrink on mastered prompts. This trend is expected and desirable: it reflects accurate prediction rather than signal loss. Unlike GRPO’s degenerate groups, these SPO samples are not discarded, they still produce well-defined gradients and contribute to learning. Notably, even under the stricter $\tau=0.02$ tolerance, SPO’s near-zero ratio remains far below GRPO’s degenerate rate, underscoring SPO’s efficient use of compute.
@@ Appendix A SPO Initialization
We show the SPO initialization procedure in Algorithm 2. In the experiments, we use $n_{0}=8$ to have a good estimation of initial baseline tracker. We ablate the setting where we use no offline estimation and rely on the online moving estimator in Section E.
@@
Practically, one may concern about the extra cost during the offline estimation of $\hat{v}_{0}$. We note that we share the offline estimation for our experiments so that people could skip this process and directly load our datasets, and there are datasets like Polaris [2] that pre-compute accuracy for Deepseek-R1-Distill-Qwen-7B [11]. The cost can be amortized across the experiments people run themselves, and we will share more (dataset, base_model) combinations to facilitate experiment efficiency.
@@ Appendix D Training and Evaluation Details
Appendix E Ablation Studies
We conduct a series of ablation studies to dissect the core components of SPO and validate our design choices. To facilitate efficient experimentation, these studies are performed under a streamlined setting compared to our main experiments. Specifically, we utilize a batch size of $256$ prompt-response pairs, and the model is updated with $4$ gradient steps for each collected batch. All ablation results are reported on the AIME 25 benchmark, using the avg@16 metric with a maximum generation length of $16,384$ tokens.
Refer to caption (a) SPO vs. $A^{*}$-PO Refer to caption (b) Baseline Ablation Refer to caption (c) Offline Initialization Ablation
Figure 6: Ablation studies evaluating the core components of SPO. (a) SPO’s adaptive baseline outperforms the static baseline of $A^{*}$-PO, demonstrating the benefit of a value function that evolves with the policy. (b) Removing the value tracker (“w/o Baseline”) causes a severe performance drop, confirming its critical role in reducing gradient variance. (c) Eliminating the offline initialization step (“w/o Offline Init”) leads to initial training instability and suboptimal convergence, highlighting the importance of a warm start for the value tracker.
SPO vs. $A^{*}$-PO. This experiment, presented in Figure 6(a), compares our proposed SPO with $A^{*}$-PO [6]. $A^{*}$-PO utilizes a static baseline derived from a pre-computed optimal value function, $V^{*}$, which is tied to the KL-regularized objective with respect to an initial reference policy, $\pi_{\text{ref}}$. While this approach is highly efficient, its central assumption may be challenged in tool-calling scenarios. In these tasks, learning involves acquiring new functional capabilities, leading to a significant policy drift where the learned policy, $\pi_{t}$, diverges substantially from $\pi_{\text{ref}}$. Consequently, the pre-computed $V^{*}$ may become a less representative baseline for the current policy’s true value function, $V_{\pi_{t}}$, potentially affecting the accuracy of the advantage estimates. In contrast, SPO’s baseline is adaptive, dynamically tracking an estimate of $V_{\pi_{t}}$ as the policy evolves. The empirical results, which show SPO’s superior performance, suggest that this adaptability is crucial. By maintaining a baseline that remains relevant to the current policy, SPO provides a more stable and effective learning signal in environments that demand significant policy evolution. Finally, from a practical perspective, $\pi_{\text{ref}}$ computation during $A^{*}$-PO policy update occupies an extra trunk of GPU memory, making it less appealing than the proposed SPO algorithm.
Baseline Ablation. Figure 6(b) presents a crucial ablation that validates the fundamental principle of using a baseline for variance reduction. In this experiment, we remove the value tracker component $\hat{v}_{-1}(x)$ from the advantage calculation, causing the algorithm to rely solely on the globally batch-normalized raw reward $r(x,y)$ as its learning signal. However, the substantial performance degradation observed is a classic illustration of the remaining challenges. While global normalization effectively controls the overall scale of rewards, the raw reward signal is still noisy on a per-sample basis as it fails to account for prompt-specific difficulty. SPO’s history-informed baseline is designed to subtract this expected difficulty, thereby effectively reducing variance and providing a much cleaner, more reliable gradient for learning. This experiment confirms that the adaptive value tracker is the most critical component for SPO’s success, directly addressing the core challenge of variance in single-stream policy optimization.
Offline Initialization Ablation. In Figure 6(c), we analyze the impact of the value tracker’s initialization phase. The standard SPO algorithm initializes the value tracker with estimates computed from a small set of $n_{0}$ offline samples, giving it a “warm start”. The ablation removes this step, forcing the tracker to learn from scratch online. The results clearly demonstrate the benefit of the offline initialization. Without it, the tracker begins with a highly inaccurate baseline, leading to high-variance gradients and significant instability in the initial training phase, as evidenced by the performance dip. Although the model eventually recovers, it fails to reach the same level of performance as the properly initialized model, underscoring the importance of a good initial value estimate for stable and effective learning.